OpenStudy (anonymous):

which arrangement for these fractions is from least to greatest 2/3 3/5 5/7

OpenStudy (anonymous):

i think its 3/5, 2/3, 3/5 is that right?

OpenStudy (texaschic101):

it might make it easier if you turned the fractions to decimals

OpenStudy (texaschic101):

2/3 = 0.666 3/5 = 0.60 5/7 = 0.71 now try it

OpenStudy (texaschic101):

crazy girl.....you got 3/5 listed twice in your answer

OpenStudy (anonymous):

Hi! You can re-name the fractions to better compare them. The denominators are 3,5,7; the least common denominator in this case is the product (because they don't have any shared factors): LCD = 3*5*7 = 105. Then, 3/5 = (multiply top & bottom by 35) = 63/105 2/3 = 70/105 5/7 = 75/105 so the correct ordering from least to greatest is: 3/5, 2/3, 5/7 Getting the decimal approximations works well, too (as another person suggested). It's great to have different tools to attack the same problem! Have a great day...

OpenStudy (anonymous):

3/5 2/3 5/7

OpenStudy (anonymous):

thanks